please give detail solution. Let X be an r.v. with uniform distribution on [0, 1]. Show that X 2 ∼ Beta(1,1).

Let X be an r.v. with uniform distribution on [0, 1]. Show that X2 ~ Beta(3, 1).
Answer #1

XNU(0,1) .: Pdf of x: f/60-1 ; 012. cdg fx : F.() = P(x+2) = ffice) de = fidze: * Let Y= x, Oxz214 Okkal y oxy</ [Note: Y=y,